OpenStudy (anonymous):

Differentiate sqrt(3x)/sin(x)

OpenStudy (anonymous):

umm i would use protuct rule, are you familiar with it?

OpenStudy (anonymous):

\[\frac{ \sqrt{3x} }{ \sin(x) }\]

OpenStudy (anonymous):

product*

OpenStudy (anonymous):

you mean quotient rule right? @HawkCrimson

OpenStudy (anonymous):

yea sry

OpenStudy (anonymous):

ok i get it now

OpenStudy (anonymous):

so id set it up like \[(3x)^{\frac{ 1 }{ 2 }}(\sin(x))^{-1}\]

OpenStudy (anonymous):

or you could just use the quotient rule

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

and use quotient rule, feel free to let me see wat you get :)

OpenStudy (anonymous):

\[\frac{ \frac{ 3 }{ 2 }x(\sin(x))-\sqrt{3x}(\cos(x)) }{ 3x }\]

OpenStudy (anonymous):

so i just simplify that: @HawkCrimson